# Berkovich Motives

### v1

@@ 1. Introduction
This paper is motivated by an application to the geometrization of the local Langlands correspondence [FS21], namely to [FS21, Conjecture I.9.5] on the independence of $\ell$ of the Langlands parameters constructed there. This application is the subject of the companion paper [Sch].
In [FS21] we used the theory of étale cohomology of rigid-analytic varieties [Hub96] and their generalization, diamonds [Sch17]. Étale cohomology only works well with torsion coefficients prime to the characteristic, naturally leading to $\ell$-adic coefficients for some auxiliary prime $\ell\neq p$. Grothendieck’s vision was that these different $\ell$-adic cohomologies should naturally be shadows of an underlying “motive”. While a full theory of motives remains elusive, much progress was made by Voevodsky et. al. through the introduction of a stable $\infty$-category $\mathcal{D}_{\mathrm{mot}}(S)$ of motivic sheaves on $S$, for any scheme $S$, equipped with all six functors [Ayo07]. Having six functors with all usual properties is sufficient for many applications, although usually one eventually runs into the issue that the categories $\mathcal{D}_{\mathrm{mot}}(S)$ remain inexplicit, even when $S$ is a geometric point. Namely, rationally morphisms in this category are governed by rational $K$-theory groups and these remain challenging to compute, being the subject of the Hodge and Tate conjecture, or also the Beilinson–Soulé vanishing conjecture. Notably, endowing $\mathcal{D}_{\mathrm{mot}}(S)$ with the “motivic $t$-structure” remains unknown.
@@
The purpose of this paper is to observe that one can also build the theory from scratch, in a language that is very close to the one used to discuss étale cohomology of diamonds [Sch17], which in particular immediately applies to arbitrary v-stacks. We decided, however, to write the results in the Berkovich setting instead of the adic setting, for the following reasons:
@@
    Archimedean Banach fields, which are by Gelfand–Mazur just $\mathbb{R}$ or $\mathbb{C}$, with norm a power of the standard norm.
@@
We sidestep the problem of defining a geometric category of Berkovich spaces and instead directly endow the (opposite of) the category of Banach rings with a very fine Grothendieck topology, and will proceed by looking at the corresponding category of stacks. The following notion is analogous to the notion of v-covers used in [Sch17], but as in Berkovich spaces one only records rank $1$ valuations it is closer to the arc-topology defined for schemes in [BM21].
@@
There are the usual set-theoretic issues that the category of Banach rings is not small. For any uncountable regular cardinal $\kappa$, one can restrict to $\kappa$-small Banach rings; and it is easy to see that when increasing $\kappa$, the corresponding pullback functors on categories of arc-sheaves are fully faithful. We will work with the union of these categories over all $\kappa$; this is the category of small arc-sheaves, i.e. those that are a small colimit of $\mathcal{M}_{\mathrm{arc}}(A)$’s. Equivalently, these are those arc-sheaves that also commute with $\kappa$-filtered colimits, for some $\kappa$.
@@
A Banach ring $A$ is strictly totally disconnected if $A$ is uniform, $\mathcal{M}(A)$ is a profinite set, and for all $x\in\mathcal{M}(A)$ the corresponding residue field $K(x)$ is algebraically closed and non-discrete.
@@
There is a natural tensor product $-\otimes_{\mathrm{mot}}-$ on $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$. More precisely, the tensor product on $\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$ restricts to a tensor product on the subcategory of finitary sheaves; and ball-invariance naturally defines a Verdier quotient $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ to which the tensor product descends. Explicitly, one has to make the arc-tensor product ball-invariant again.
@@
where $\mathcal{O}_{<1}\subset\mathcal{O}$ is the sheaf of elements of norm $<1$. Then the natural map
@@
    where $K(C_{<1})$ denotes the algebraic $K$-theory of the non-unital, but Tor-unital, ring $C_{<1}\subset C$.
@@
$-\otimes_{\mathrm{mot}}\mathbb{Z}(1):\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)\to\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$
@@
Remark 1.12.
The proof we give is substantially simpler than the proof of Ayoub–Gallauer–Vezzani [AGV22]. We do not need to use formal models.
@@
Here, we use the notion of rigidity of Gaitsgory–Rozenblyum [GR17, Definition 9.1.2]; it is an adaptation to the non-compactly generated case of the idea that the compact and dualizable objects agree. Again, this rigidity in the relative case is something that does not usually hold true, not even for étale sheaves on schemes or rigid-analytic varieties. Namely, in those settings the category is compactly generated, but not all compact objects are dualizable (as there are constructible sheaves that are not locally constant). The Berkovich setting yields categories closer to the category of sheaves on the compact Hausdorff space $\mathcal{M}(A)$, and the latter is rigid.
@@
Acknowledgments. The idea of using rigid-analytic motives to study independence of $\ell$ of $L$-parameters was found while giving the lecture course on $6$-functor formalisms in the winter term 2022/23. This was of course inspired by the work of Richarz and Scholbach on motivic geometric Satake, and their work in the direction of a motivic version of the results of V. Lafforgue. I want to thank Johannes Anschütz, Joseph Ayoub, Dustin Clausen and David Hansen for very helpful feedback on a draft. The results of this paper were the basis for an ARGOS seminar in the winter term 2024/25, and I heartily thank all the speakers and participants for their feedback.
@@ 2. Banach rings and Berkovich spectrum
For the case of seminormed rings, it suffices to handle the initial object, and the case of filtered colimits and of pushouts. The initial object is given by $\mathbb{Z}$ with its standard norm $|n|_{\mathbb{Z}}=\pm n$. (Here we use the condition |-$|\text{-}1|_{R}\leq 1$ on a seminorm, which implies $|1|_{R}\leq 1$ and then the bound follows from the triangle inequality.) Filtered colimits $\mathrm{colim}_{i}R_{i}$ are computed naively on the underlying rings, and the norm of an element $r\in\mathrm{colim}_{i}R_{i}$ that is the image of some $r_{i_{0}}\in R_{i_{0}}$ is given as the infimum of $|r_{i}|_{R_{i}}$, where $r_{i}$ is the image of $r_{i_{0}}$ in $R_{i}$ for $i\geq i_{0}$. For pushouts, take a diagram $B\leftarrow A\to C$ of seminormed rings. On $D=B\otimes_{A}C$, we can define the projective norm: Namely, we set
@@ 2.1. Free seminormed rings
    A morphism $f:X\to Y$ is compact if for any filtered colimit $Z=\mathrm{colim}_{i}Z_{i}$ with a map $Y\to Z$, there is some $i$ and a map $X\to Z_{i}$ making
@@
    The $\infty$-category is compactly assembled if it is generated under colimits by objects of the form
@@
    where all maps $X_{i}\to X_{i+1}$ are compact.
@@
has a left adjoint.
@@
It follows that $\mathbb{Z}[T]_{r}=\mathrm{colim}_{r^{\prime}>r}\mathbb{Z}[T]_{r^{\prime}}$ is a presentation as a sequential colimit along compact transition maps. The whole category is generated by this example (for varying $r\in\mathbb{R}_{>0}$) under colimits. ∎
@@ 2.2. Banach fields
There are three flavours of Banach fields: The discrete Banach fields in which $K$ has the discrete topology and $|\cdot|_{K}$ takes values in $\{0,1\}$, and the nondiscrete fields in which there exists some $\pi\in K$ with $10<|\pi|_{K}<$ 1, which can further be divided into the archimedean and nonarchimedean ones. In the following, by a nonarchimedean Banach field we also mean that in particular it is not discrete.
@@
We are free to replace $R$ by its completion. We can also replace $R$ by the quotient by any proper closed ideal. Note that any maximal ideal $\mathfrak{m}$ of $R$ is already closed. Otherwise, its closure is all of $R$, and in particular contains $1$. Thus, there exists $x\in\mathfrak{m}$ with $|1-x|<1$. Thus, $x=1-y$ with $|y|_{R}<1$. But then $1+y+y^{2}+\ldots$ converges in $R$ and defines $(1-y)^{-1}$. Thus, $x\in\mathfrak{m}$ is invertible, which is a contradiction.
In other words, we can assume that $R=K$ is a field. If we consider all possible seminorms on $K$ bounded by the given one and with $|1|_{K}=1$, then this category forms a partially ordered set and has all filtered colimits; we can thus pick a maximal element. We can also pass to the completion. Our task now is to prove that $K$ is a Banach field. First note that $x\mapsto|x^{2}|_{K}^{1/2}$ defines another norm on $K$ which is smaller; by choice of $K$, we thus have $|x|_{K}^{2}=|x^{2}|_{K}$ and more generally $|x|_{K}^{n}=|x^{n}|_{K}$ for $n>0$. Now for any nonzero $x\in K$ we can define a new norm by sending $y$ to the infimum over all $n>0$ of $|x|_{K}^{-n}|x^{n}y|_{K}$. This still sends $1\in K$ to $1\in\mathbb{R}_{\geq 0}$ as $|x|_{K}^{n}=|x^{n}|_{K}$. Again, the choice of $K$ implies that we must thus have $|x|_{K}^{-n}|x^{n}y|_{K}=|y|_{K}$, and in particular $|xy|_{K}=|x|_{K}|y|_{K}$. ∎
@@
Let $R$ be a Banach ring and $f\in R$. Then $f$ is invertible if and only if $||f||:\mathcal{M}(R)\to\mathbb{R}_{\geq 0}$ is everywhere strictly positive.
@@
If $f$ is not invertible, then the closure of the ideal generated by $f$ is not all of $R$, and hence the quotient $S$ is a nonzero Banach ring where $f$ maps to 00. Any point in the image of $\mathcal{M}(S)\to\mathcal{M}(R)$ is then a point where $||f||$ is zero. ∎
The following notion was introduced by Kedlaya–Liu [KL15] under the name “free of trivial spectrum”; later, Kedlaya introduced the term “analytic”. While this clashes with the notion of “analytic ring” in condensed mathematics, in the present paper it does not lead to problems.
@@
It is not possible to define a sheaf of Banach rings on $\mathcal{M}(R)$ localizing $R$. Using condensed mathematics (and in particular the discussion in [CS22]), one can still produce a sheaf of animated gaseous $R$-algebras, at least when $R$ is analytic (which can always be ensured by passing to $R[T^{\pm 1}]_{1/2}$). We will only use the following seemingly elementary fact, which can be proved directly.
@@
The previous corollary implies injectivity of the map from idempotent elements to open and closed subsets. The high-tech proof of the converse is the following. We have to see that any open and closed decomposition $\mathcal{M}(R)=U_{1}\sqcup U_{2}$ yields an idempotent element in $R$. Passing to the completion of $R[T^{\pm 1}]_{1/2}$ we may assume that $R$ is analytic (the idempotent will necessarily only have a constant term, as it is invariant under rescaling the variable $T$). Then we have a sheaf $\mathcal{O}$ of animated gaseous $R$-algebras on $\mathcal{M}(R)$, and the open and closed decomposition yields a global section $e\in R$ that is 00 on one part and $1$ on the other.
A direct proof goes as follows; this uses the notion of uniform completions and the Gelfand transform introduced just below. Assume $T\in R$ and $r>0$. We have maps $R\to R\langle T\rangle_{r}$ and $R\to R\langle T^{-1}\rangle_{r}$. We claim that if one has idempotents $e_{1}\in R\langle T\rangle_{r}$ and $e_{2}\in R\langle T^{-1}\rangle_{r}$ that map to the same idempotent in $R\langle T^{\pm 1}\rangle_{r}$, then they come from a unique idempotent $e\in R$. Once one knows this, the result follows by a standard localization argument. One first checks that idempotents lift uniquely from the uniform completion of $R$ back to $R$: Given any sufficiently good approximation, the Newton method will produce an idempotent (as the equation $x^{2}-x=0$ is finite étale). Thus, we can assume that $R$ is uniform. By a similar argument, we can also replace $R\langle T\rangle_{r}$ by the colimit of $R\langle T\rangle_{r^{\prime}}$ over $r^{\prime}>r$; we denote this by $R\langle T\rangle_{r}^{\dagger}$. Then the map
$R\to\mathrm{eq}(R\langle T\rangle_{r}^{\dagger}\times R\langle T^{-1}\rangle_{r}^{\dagger}\rightrightarrows R\langle T^{\pm 1}\rangle_{r}^{\dagger})$
is actually an isomorphism. Injectivity follows from the Gelfand transform being injective, and surjectivity from an argument of Kedlaya–Liu [KL15]. More precisely, the main point is that
$R\langle T\rangle_{r}^{\dagger}=R\langle U\rangle_{r}^{\dagger}/(T-U)$
where $U$ is a free variable, and the ideal generated by $T-U$ is closed. When $R$ is uniform, this assertion can be reduced to the case of Banach fields, where it is standard. The same applies to the other two overconvergent algebras, and then the result follows by a direct diagram chase. ∎

### v2

@@ 1. Introduction
This paper is motivated by an application to the geometrization of the local Langlands correspondence [FS21], namely to [FS21, Conjecture I.9.5] on the independence of $\ell$ of the Langlands parameters constructed there. This application is the subject of the companion paper [Sch25a].
In [FS21] we used the theory of étale cohomology of rigid-analytic varieties [Hub96] and their generalization, diamonds [Sch17]. Étale cohomology only works well with torsion coefficients prime to the characteristic, naturally leading to $\ell$-adic coefficients for some auxiliary prime $\ell\neq p$. Grothendieck’s vision was that these different $\ell$-adic cohomologies should naturally be shadows of an underlying “motive”. While a full theory of motives remains elusive, much progress was made by Voevodsky et. al. through the introduction of a stable $\infty$-category $\mathcal{D}_{\mathrm{mot}}(S)$ of motivic sheaves on $S$, for any scheme $S$, equipped with all six functors [Ayo07]. Having six functors with all usual properties is sufficient for many applications, although usually one eventually runs into the issue that the categories $\mathcal{D}_{\mathrm{mot}}(S)$ remain inexplicit, even when $S$ is a geometric point. Namely, rationally morphisms in this category are governed by rational $K$-theory groups and these remain challenging to compute, being the subject of the Hodge and Tate conjectures, and also the Beilinson–Soulé vanishing conjecture. Notably, endowing $\mathcal{D}_{\mathrm{mot}}(S)$ with the “motivic $t$-structure” remains unknown.
@@
The purpose of this paper is to observe that one can also build the theory from scratch, in a language that is very close to the one used to discuss étale cohomology of diamonds [Sch17] – in particular, making use of very fine Grothendieck topologies like the v-topology, and strictly totally disconnected spaces. We decided, however, to write the results in the Berkovich setting instead of the adic setting, for the following reasons:
@@
    Archimedean Banach fields, which are by Gelfand–Mazur just $\mathbb{R}$ or $\mathbb{C}$, with norm a power $|\cdot|^{\alpha}$, $\alpha\in(0,1]$, of the standard norm $|\cdot|$.
@@
We sidestep the problem of defining a geometric category of Berkovich spaces and instead directly endow (the opposite of) the category of Banach rings with a very fine Grothendieck topology, and will proceed by looking at the corresponding category of stacks. The following notion is analogous to the notion of v-covers used in [Sch17], but as in Berkovich spaces one only records rank $1$ valuations it is closer to the arc-topology defined for schemes in [BM21].
@@
There are the usual set-theoretic issues that the category of Banach rings is not small. For any uncountable regular cardinal $\kappa$, one can restrict to $\kappa$-small Banach rings, meaning those whose underlying set has cardinality less than $\kappa$; and it is easy to see that when increasing $\kappa$, the corresponding pullback functors on categories of arc-sheaves are fully faithful. We will work with the union of these categories over all $\kappa$; this is the category of small arc-sheaves, i.e. those that are a small colimit of $\mathcal{M}_{\mathrm{arc}}(A)$’s. Equivalently, these are those arc-sheaves that also commute with $\kappa$-filtered colimits, for some $\kappa$.
@@
A Banach ring $A$ is strictly totally disconnected if $A$ is uniform [^1] The notion of uniformity is recalled in Definition 2.20., $\mathcal{M}(A)$ is a profinite set, and for all $x\in\mathcal{M}(A)$ the corresponding residue field $K(x)$ is algebraically closed and non-discrete.
@@
There is a natural tensor product $\otimes_{\mathrm{mot}}-$ on $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$. More precisely, the tensor product on $\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$ restricts to a tensor product on the subcategory of finitary sheaves; and ball-invariance naturally defines a Verdier quotient $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ to which the tensor product descends. Explicitly, one has to make the arc-tensor product ball-invariant again.
@@
where $\mathcal{O}_{<1}\subset\mathcal{O}$ is the sheaf of elements of norm <1<1. Then the natural map
@@
    where $K(C_{<1})$ denotes the algebraic $K$-theory of the non-unital, but Tor-unital, ring $C_{<1}\subset C$. [^2] Recall that a non-unital ring $A$ is Tor-unital if its unitalization $A^{+}=A\oplus\mathbb{Z}$ satisfies $\mathbb{Z}\otimes^{\mathbb{L}}_{A^{+}}\mathbb{Z}=\mathbb{Z}$. For Tor-unital rings, one defines $K(A)$ as the fibre of $K(A^{+})\to K(\mathbb{Z})$, and it is a theorem of Suslin [Sus95] that this agrees with the fibre of $K(A^{\prime})\to K(A^{\prime}/A)$ for any unital ring $A^{\prime}$ containing $A$ as an ideal.
@@
$\otimes_{\mathrm{mot}}\mathbb{Z}(1):\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)\to\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$
@@
Here, we use the notion of rigidity of Gaitsgory–Rozenblyum [GR17, Chapter I, Definition 9.1.2]; it is an adaptation to the non-compactly generated case of the idea that the compact and dualizable objects agree. Again, this rigidity in the relative case is something that does not usually hold true, not even for étale sheaves on schemes or rigid-analytic varieties. Namely, in those settings the category is compactly generated, but not all compact objects are dualizable (as there are constructible sheaves that are not locally constant). The Berkovich setting yields categories closer to the category of sheaves on the compact Hausdorff space $\mathcal{M}(A)$, and the latter is rigid.
@@
Acknowledgments. The idea of using rigid-analytic motives to study independence of $\ell$ of $L$-parameters was found while giving the lecture course on $6$-functor formalisms in the winter term 2022/23. This was of course inspired by the work of Richarz and Scholbach on motivic geometric Satake, and their work in the direction of a motivic version of the results of V. Lafforgue. I want to thank Johannes Anschütz, Joseph Ayoub, Dustin Clausen, David Hansen, Hyungseop Kim and Dmitry Kubrak for very helpful feedback on a draft, and Mura Yakerson for some discussions on the cancellation theorem. The results of this paper were the basis for an ARGOS seminar in the winter term 2024/25, and I heartily thank all the speakers and participants for their feedback. Moreover, I am very grateful to the referee for a very careful reading, and a large number of suggestions for improvements.
@@ 2. Banach rings and Berkovich spectrum
For the case of seminormed rings, it suffices to handle the initial object, and the case of filtered colimits and of pushouts. The initial object is given by $\mathbb{Z}$ with its standard norm $|n|_{\mathbb{Z}}=\pm n$. (Here we use the condition $|\text{-}1|_{R}\leq 1$ on a seminorm, which implies $|1|_{R}\leq 1$ and then the bound follows from the triangle inequality.) Filtered colimits $\mathrm{colim}_{i}R_{i}$ are computed naively on the underlying rings, and the norm of an element $r\in\mathrm{colim}_{i}R_{i}$ that is the image of some $r_{i_{0}}\in R_{i_{0}}$ is given as the infimum of $|r_{i}|_{R_{i}}$, where $r_{i}$ is the image of $r_{i_{0}}$ in $R_{i}$ for $i\geq i_{0}$. For pushouts, take a diagram $B\leftarrow A\to C$ of seminormed rings. On $D=B\otimes_{A}C$, we can define the projective norm: Namely, we set
@@ 2.1. Free seminormed rings
    A morphism $f:X\to Y$ is weakly compact if for any filtered colimit $Z=\mathrm{colim}_{i}Z_{i}$ with a map $Y\to Z$, there is some $i$ and a map $X\to Z_{i}$ making
@@
    The $\infty$-category $C$ is compactly assembled if filtered colimits commute with finite limits in $C$, and it is generated under colimits by objects of the form
@@
    where all maps $X_{i}\to X_{i+1}$ are weakly compact.
@@
has a left adjoint, see [Ram24a, Theorem 2.39]. The terminology “weakly compact” follows [Ram24a, Definition 2.1]. As it is the only notion of compactness of morphisms we will use, we will occasionally drop the “weakly”. As in a compactly assembled $\infty$-category, the different notions of compact morphisms notions agree (cf. [Ram24a, Example 2.20]), we hope this will not lead to confusion.
@@
It follows that $\mathbb{Z}[T]_{r}=\mathrm{colim}_{r^{\prime}>r}\mathbb{Z}[T]_{r^{\prime}}$ is a presentation as a sequential colimit along weakly compact transition maps. The whole category is generated by this example (for varying $r\in\mathbb{R}_{>0}$) under colimits. ∎
@@ 2.2. Banach fields
There are three flavours of Banach fields: The discrete Banach fields in which $K$ has the discrete topology and $|\cdot|_{K}$ takes values in $\{0,1\}$, and the nondiscrete fields in which there exists some $\pi\in K$ with $0<|\pi|_{K}<1$, which can further be divided into the archimedean and nonarchimedean ones. In the following, by a nonarchimedean Banach field we also mean that in particular it is not discrete.
@@
We are free to replace $R$ by its completion. We can also replace $R$ by the quotient by any proper closed ideal. Note that any maximal ideal $\mathfrak{m}$ of $R$ is already closed. Otherwise, its closure is all of $R$, and in particular contains $1$. Thus, there exists $x\in\mathfrak{m}$ with $|1-x|_{R}<1$. Thus, x=1yx=1-y with $|y|_{R}<1$. But then $1+y+y^{2}+\ldots$ converges in $R$ and defines $(1-y)^{-1}$. Thus, $x\in\mathfrak{m}$ is invertible, which is a contradiction.
In other words, we can assume that $R=K$ is a field. If we consider all possible seminorms on $K$ bounded by the given one and with $|1|_{K}=1$, then this category forms a partially ordered set (where maps are morphisms of $K$-Banach rings) and has all filtered colimits; we can thus pick a maximal element. We can also pass to the completion. Our task now is to prove that $K$ is a Banach field. First note that $x\mapsto\mathrm{lim}_{n\to\infty}|x^{n}|_{K}^{1/n}$ defines another norm on $K$ which is smaller; by choice of $K$, we thus have $|x|_{K}^{n}=|x^{n}|_{K}$ for $n>0$. Now for any nonzero $x\in K$ we can define a new norm by sending $y$ to the infimum over all $n>0$ of $|x|_{K}^{-n}|x^{n}y|_{K}$. This still sends $1\in K$ to $1\in\mathbb{R}_{\geq 0}$ as $|x|_{K}^{n}=|x^{n}|_{K}$. Again, the choice of $K$ implies that we must thus have $|x|_{K}^{-n}|x^{n}y|_{K}=|y|_{K}$, and in particular $|xy|_{K}=|x|_{K}|y|_{K}$. ∎
@@
Let $R$ be a Banach ring and $f\in R$. Then $f$ is invertible if and only if $f||:\mathcal{M}(R)\to\mathbb{R}_{\geq 0}$ is everywhere strictly positive.
@@
If $f$ is not invertible, then the closure of the ideal generated by $f$ is not all of $R$ (as maximal ideals are closed, as established in the previous proof), and hence the quotient $S$ is a nonzero Banach ring where $f$ maps to 0. Any point in the image of $\mathcal{M}(S)\to\mathcal{M}(R)$ is then a point where $f||$ is zero. ∎
The following notion was introduced by Kedlaya–Liu [KL15, Remark 2.3.9 (d)] under the name “free of trivial spectrum”; later, Kedlaya [BCKW19, Definition 1.1.2] introduced the term “analytic”. While this clashes with the notion of “analytic ring” in condensed mathematics, in the present paper it does not lead to problems.
@@
It is not possible to define a sheaf of Banach rings on $\mathcal{M}(R)$ localizing $R$. Using condensed mathematics (and in particular the discussion in [CS22, Lecture 5, Corollary 5.10] generalized from $\mathbb{C}$-algebras to general analytic Banach algebras), one can still produce a sheaf of animated gaseous $R$-algebras, at least when $R$ is analytic (which can always be ensured by passing to $R[T^{\pm 1}]_{1/2}$). We will only use the following seemingly elementary fact, which can be proved directly.
@@
The previous corollary implies injectivity of the map from idempotent elements to open and closed subsets. The high-tech proof of the converse is the following. We have to see that any open and closed decomposition $\mathcal{M}(R)=U_{1}\sqcup U_{2}$ yields an idempotent element in $R$. Passing to the completion of $R[T^{\pm 1}]_{1/2}$ we may assume that $R$ is analytic (the idempotent will necessarily only have a constant term, as it is invariant under rescaling the variable $T$). Then we have a sheaf $\mathcal{O}$ of animated gaseous $R$-algebras on $\mathcal{M}(R)$, and the open and closed decomposition yields a global section $e\in R$ that is 0 on one part and $1$ on the other.
A direct proof goes as follows; this uses the notion of uniform completions and the Gelfand transform introduced just below. Assume $f\in R$ and $r>0$. We have maps $R\to R\langle F\rangle_{r}$ and $R\to R\langle f^{-1}\rangle_{r}$, whose Berkovich spaces are given by the subsets $\{|f|\leq r\}$ and $\{|f|\geq r\}$ of $\mathcal{M}(R)$. We claim that if one has idempotents $e_{1}\in R\langle f\rangle_{r}$ and $e_{2}\in R\langle f^{-1}\rangle_{r}$ that map to the same idempotent in $R\langle f^{\pm 1}\rangle_{r}$, then they come from a unique idempotent $e\in R$. Once one knows this, the result follows by a standard localization argument.
One first checks that idempotents lift uniquely from the uniform completion of $R$ back to $R$: Given any sufficiently good approximation, the Newton method will produce an idempotent (as the equation $x^{2}-x=0$ is finite étale). Thus, we can assume that $R$ is uniform. By a similar argument, we can also replace $R\langle f\rangle_{r}$ by the colimit of $R\langle f\rangle_{r^{\prime}}$ over $r^{\prime}>r$; we denote this by $R\langle f\rangle_{r}^{\dagger}$. Then the map
$R\to\mathrm{eq}(R\langle f\rangle_{r}^{\dagger}\times R\langle f^{-1}\rangle_{r}^{\dagger}\rightrightarrows R\langle f^{\pm 1}\rangle_{r}^{\dagger})$
is actually an isomorphism, and in particular the gluing of idempotents follows. Injectivity follows from the Gelfand transform being injective, and surjectivity from an argument of Kedlaya–Liu [KL15, Corollary 2.8.9]. More precisely, the main point is that
$R\langle f\rangle_{r}^{\dagger}=R\langle U\rangle_{r}^{\dagger}/(U-f)$
where $U$ is a free variable, and the ideal generated by UfU-f is closed. When $R$ is uniform, this assertion can be reduced to the case of Banach fields, where it is standard. The same applies to the other two overconvergent algebras, and then the result follows by a direct diagram chase. ∎